# Chenghao Xiao  G Thomas Hudson Noura Al Moubayed

### v1

@@ 1 Introduction
Despite the promise shown by dense retrievers in fetching references for knowledge-intensive tasks, these systems still fall short in retrieving reliable and cite-worthy references (Malaviya et al., 2023), compared to state-of-the-art proprietary LLMs (OpenAI, 2023) in a standalone manner, highlighting the undesirable behavior of retrievers in assisting LLMs. This discrepancy is more pronounced in reasoning-intensive tasks, where retrieval-augmented generation methods present inconsistent gains, or even performance degradation to LLMs (Bai et al., 2023; Xu et al., 2023).
@@ 2.2 Datasets
Task (normal-$\downarrow$) Domain (normal-$\downarrow$) Dataset (normal-$\downarrow$) #MCR #Pairs #Query #Query #Corpusnormal-$\clubsuit$ Query Document
@@ 2.3 Models
We first benchmark the performance of state-of-the-art bi-encoders, spanning across unsupervised dense retrieval models (U-DR), supervised dense retrieval models (S-DR), supervised instruction-aware dense retrieval models (I-DR). For U-DR models, we have Contriever (Izacard et al., 2022). For S-DR models, we include two Sentence Transformers (Reimers & Gurevych, 2019) models that have been the most popular in the last few years, all-mpnet-base-v2 and all-MiniLM-L6-v2; and Dragon+ (Lin et al., 2023), a state-of-the-art S-DR model that is progressively distilled with different views of other SOTA models. For I-DR models, we include InstructOR (Su et al., 2022), BGE (Xiao et al., 2023c) and BGE-M3 (Chen et al., 2024), and consider all of their model sizes (base, large, XL for InstructOR and small, base, large for BGE) to understand the scaling laws (if there’s any) between model size and reasoning abilities. We further include E5-Mistral (Wang et al., 2023a), a latest state-of-the-art instruction-aware embedding model that is finetuned from Mistral (Jiang et al., 2023) with {instruction, query, answer} pairs distilled from GPT-4. We recognize the recent advancements in unifying generative and representation abilities into one model, and benchmark Grit (Muennighoff et al., 2024), whose data and training process of the embedding abilities are similar to Jiang et al. (2023), but differing in its full parameters finetuning and joint learning with the generative objective. Lastly, we benchmark popular proprietary models Cohere Embed-English-v3.0 and OpenAI text-embedding-ada-002, text-embedding-3-small, text-embedding-3-large through API encoding.
@@ 3.1 Full-dataset Retrieval Setting
Contriever w/o Inst. 0.318 0.144 0.246 0.013 0.050 0.086 0.471 0.108 0.109 0.308 0.093 0.126
w/ Inst. 0.271 0.177 0.217 0.009 0.049 0.076 0.263 0.105 0.106 0.218 0.071 0.106
all-mpnet-base-v2 w/o Inst. 0.224 0.263 0.290 0.024 0.034 0.118 0.207 0.058 0.002 0.718 0.531 0.110
w/ Inst. 0.020 0.130 0.272 0.013 0.030 0.104 0.097 0.044 0.010 0.692 0.488 0.079
@@
Dragon+ w/o Inst. 0.321 0.277 0.280 0.020 0.041 0.089 0.672 0.087 0.103 0.451 0.176 0.151
w/ Inst. 0.252 0.241 0.264 0.017 0.042 0.082 0.609 0.081 0.108 0.362 0.128 0.134
InstructOR-base w/o Inst. 0.246 0.263 0.282 0.024 0.036 0.102 0.301 0.059 0.036 0.584 0.415 0.135
w/ Inst. 0.201 0.239 0.254 0.023 0.038 0.096 0.164 0.052 0.069 0.582 0.401 0.128
InstructOR-large w/o Inst. 0.248 0.295 0.332 0.037 0.051 0.125 0.264 0.068 0.016 0.710 0.546 0.147
w/ Inst. 0.234 0.266 0.286 0.031 0.044 0.108 0.226 0.062 0.034 0.672 0.527 0.142
@@
Cohere-embedding-v3 w/o Inst. 0.151 0.263 0.285 0.043 0.041 0.099 0.580 0.151 0.038 0.723 0.572 0.167
w/ Inst. 0.187 0.290 0.279 0.050 0.078 0.101 0.650 0.171 0.033 0.721 0.566 0.187
OpenAI-Ada-002 w/o Inst. 0.256 0.293 0.310 0.031 0.058 0.133 0.197 0.100 0.042 0.732 0.834 0.170
w/ Inst. 0.106 0.248 0.239 0.026 0.058 0.118 0.114 0.096 0.048 0.673 0.824 0.139
OpenAI-3-small w/o Inst. 0.306 0.309 0.337 0.030 0.061 0.146 0.315 0.125 0.066 0.711 0.720 0.191
w/ Inst. 0.212 0.272 0.296 0.030 0.066 0.138 0.255 0.129 0.036 0.643 0.721 0.166
OpenAI-3-large w/o Inst. 0.373 0.341 0.420 0.034 0.101 0.240 0.291 0.164 0.075 0.901 0.896 0.237
w/ Inst. 0.342 0.314 0.375 0.050 0.136 0.212 0.339 0.211 0.074 0.877 0.894 0.252
@@
Latest instruction-aware decoder representational models (Wang et al., 2023a; Muennighoff et al., 2024) seem to start bridging this behavioral gap, pointing the promising direction of scaling decoder models for next-level representational abilities. Cohere’s proprietary Embed-English-v3.0 presents a strong gain brought by instructions. Considering its relatively early release time (Nov 02, 2023) compared to latest models that yield better absolute performance (OpenAI-v3, E5-Mistral, Grit) on RAR-b, this behavior is surprising, and might be relevant to the efforts put in relevance feedback from LLMs  [^3] https://txt.cohere.com/introducing-embed-v3/. However, no public information is available whether Cohere’s current embedding model is an encoder or a decoder, and thus its strong instruction-following abilities can not add to our main narrative of the potentail for decoder representation models.
@@ 3.3 Multiple-choice Retrieval Setting
A special case in the models is InstructOR, which is known to have seen the formats of the first 6 tasks through finetuning on Super-NI (Wang et al., 2022b), enabling it to achieve a better performance on these tasks compared to traditional S-DR models. Yet, its performance on the two settings of CSTS that we construct is stuck at chance level. Without solid control experiments, we hypothesize that this pattern stems from the subpar generalizablity of encoder models. Although the pattern can currently be empirically supported by the better performance achieved by similarly instruction-tuned decoder models including E5-Mistral, Grit and OpenAI models, the generalization gap between encoders and decoders has not yet been systematically investigated on information retrieval, and we position this as a valuable research topic.
@@ 4.1 Behavioral Analysis through the lens of Temporal Reasoning
Lastly, for the Context+Fact setting, we further concatenate the context (the background knowledge of the entity in the question that is going to be asked) after the facts. On top of the Fact setting, the Context+Fact setting further reflects a model’s robustness towards irrelevant information (since the answer can be extracted only from facts, but not from contexts), and encoding capabilities for texts of varied length ranges (Xiao et al., 2023a). However, there potentially is unfair comparison in here because of the max sequence length each model is capable of encoding. For instance, OpenAI embedidng models are able to encode sequence up to 8192 tokens and are faced with more noise to be robust against. On the other hand, models with max sequence length of 512 need to deal with less noise. Therefore, the Context-Fact setting here is only for analysis purposes, and we do not average the scores from this setting into the average TR scores in the main table. For E5-Mistral and Grit, we set its max sequence length to 8192 here to match with OpenAI models, for attaining a comparable understanding on the Context+Fact setting.
@@
Cohere-embedding-v3 w/o Inst. 0.015 0.019 0.359 0.197 0.085 0.275 0.162 0.159 0.151
@@ Appendix A Dataset Format
Dataset
Instruction
Query
Doc
$\alpha$NLI
Given the following start and end of a story, retrieve a possible reason that leads to the end.
Start: Ron started his new job as a landscaper today. End: Ron is immediately fired for insubordination.
Ron ignores his bosses’s orders and called him an idiot.
HellaSwag
Given the following unfinished context, retrieve the most plausible ending to finish it.
A man is sitting on a roof. He
starts pulling up roofing on a roof.
PIQA
Given the following goal, retrieve a possible solution.
How do I ready a guinea pig cage for it’s new occupants?
Provide the guinea pig with a cage full of a few inches of bedding made of ripped paper strips, you will also need to supply it with a water bottle and a food dish.
SocialiQA
Given the following context and question, retrieve the correct answer.
Context: Tracy didn’t go home that evening and resisted Riley’s attacks. Question: What does Tracy need to do before this?
Find somewhere to go
Quail
Given the following context and question, retrieve the correct answer.
@@
about 10 minutes
ARC-Challenge
Retrieve the answer to the question.
An astronomer observes that a planet rotates faster after a meteorite impact. Which is the most likely effect of this increase in rotation?”
Planetary days will become shorter.
Winogrande
Given the following sentence, retrieve an appropriate answer to fill in the missing underscored part.
Sentence: Sarah was a much better surgeon than Maria so _ always got the easier cases..
Maria
TempReason-L1
Given the following question about time, retrieve the correct answer.
What is the time 6 year and 4 month after Nov, 1185
Sun Mar 01 1192 00:00:00 GMT-0001 (Greenwich Mean Time)
TempReason-L2 pure
Given the following question, retrieve the correct answer.
Which employer did Jaroslav Pelikan work for in Jan, 1948?
Valparaiso University
TempReason-L2 fact
Given the following question and facts, retrieve the correct answer.
@@
Valparaiso University
TempReason-L2 context
Given the following question, facts and contexts, retrieve the correct answer.
@@
Valparaiso University
TempReason-L3 pure
Given the following question, retrieve the correct answer.
Which employer did Jaroslav Pelikan work for before Concordia Seminary?
Valparaiso University
TempReason-L3 fact
Given the following question and facts, retrieve the correct answer.
@@
Valparaiso University
TempReason-L3 context
Given the following question, facts and contexts, retrieve the correct answer.
@@
Valparaiso University
SpartQA
Given the following spatial reasoning question, retrieve the right answer.
We have three blocks. Lets call them A, B and C. Block B is below A. Block A is below C. Block A contains a medium yellow square. Block B has two medium blue squares. Medium blue square number one is touching the bottom edge of this block. Medium blue square number two is below a medium yellow square. Medium blue square number one is below the square which is below the medium yellow square. It is below the medium yellow square. Block C contains one medium black square. What is below the black thing? a medium yellow square that is in block A or a medium yellow square that is in block B?
both of them
Math-pooled
Retrieve the answer for the following math problem.
{Omitted for brevity} {Omitted for brevity}
HumanEvalPack-MBPP
Retrieve the answer for the following coding problem.
Finish the following code based on the docstring: {Omitted for brevity}
{Omitted for brevity}
CSTS-easy
Retrieve an aspect and a sentence which are similar to the following sentence.
3 people wearing festive costumes and feathers dancing in a parade .
In terms of ”The number of persons” retrieve a sentence similar to the following. Three uniquely dressed women dancing in a parade .
CSTS-hard
Retrieve a condition under which the following two sentences are similar.
Sentence1: 3 people wearing festive costumes and feathers dancing in a parade .; Sentence2: Three uniquely dressed women dancing in a parade .
The number of persons.
@@ Appendix D Full setting Recall@10
Cohere-embedding-v3 w/o Inst. 0.223 0.400 0.423 0.073 0.082 0.180 0.886 0.275 0.080 0.790 0.739 0.266

### v2

@@ 1 Introduction
Despite the promise shown by dense retrievers in fetching references for knowledge-intensive tasks, these systems still fall short in retrieving reliable and cite-worthy references (Malaviya et al., 2023), compared to state-of-the-art proprietary LLMs (OpenAI, 2023) in a standalone manner, highlighting the undesirable behavior of retrievers in assisting LLMs (BehnamGhader et al., 2023; Asai et al., 2024). This discrepancy is more pronounced in reasoning-intensive tasks, where retrieval-augmented generation methods present inconsistent gains, or even performance degradation to LLMs (Bai et al., 2023; Xu et al., 2023).
@@ 2.2 Datasets
[…]
@@ 2.3 Models
We first benchmark the performance of state-of-the-art bi-encoders, spanning across unsupervised dense retrieval models (U-DR), supervised dense retrieval models (S-DR), supervised instruction-aware dense retrieval models (I-DR). For U-DR models, we have Contriever (Izacard et al., 2022). For S-DR models, we include two Sentence Transformers (Reimers & Gurevych, 2019) models that have been the most popular in the last few years, all-mpnet-base-v2 and all-MiniLM-L6-v2; Dragon+ (Lin et al., 2023), an S-DR model that is progressively distilled with different views of other SOTA models, and BGE-M3 (Chen et al., 2024). For I-DR models, we include Instructor (Su et al., 2022), BGE (v1.5) (Xiao et al., 2023c), and consider all of their model sizes (base, large, XL for Instructor and small, base, large for BGE) to understand the scaling laws (if there’s any) between model size and reasoning abilities. We further include E5-Mistral (Wang et al., 2023a), a latest state-of-the-art instruction-aware embedding model that is finetuned from Mistral (Jiang et al., 2023) with {instruction, query, answer} pairs distilled from GPT-4. We recognize the recent advancements in unifying generative and representation abilities into one model, and benchmark GritLM (Muennighoff et al., 2024), whose data and training process of the embedding abilities are similar to Jiang et al. (2023), but differing in its full parameters finetuning and joint learning with the generative objective. Lastly, we benchmark popular proprietary models Cohere Embed-English-v3.0 and OpenAI text-embedding-ada-002, text-embedding-3-small, text-embedding-3-large through API encoding.
@@ 3.1 Full-dataset Retrieval Setting
Contriever w/o Inst. 0.318 0.144 0.246 0.013 0.050 0.086 0.471 0.108 0.109 0.308 0.093 0.123
w/ Inst. 0.271 0.177 0.217 0.009 0.049 0.076 0.263 0.105 0.106 0.218 0.071 0.104
all-mpnet-base-v2 w/o Inst. 0.224 0.263 0.290 0.024 0.034 0.118 0.207 0.058 0.002 0.718 0.531 0.108
w/ Inst. 0.020 0.130 0.272 0.013 0.030 0.104 0.097 0.044 0.010 0.692 0.488 0.077
@@
Dragon+ w/o Inst. 0.321 0.277 0.280 0.020 0.041 0.089 0.672 0.087 0.103 0.451 0.176 0.150
w/ Inst. 0.252 0.241 0.264 0.017 0.042 0.082 0.609 0.081 0.108 0.362 0.128 0.133
bge-m3 w/o Inst. 0.247 0.257 0.229 0.049 0.075 0.090 0.417 0.140 0.075 0.692 0.388 0.178
w/ Inst. 0.247 0.255 0.190 0.048 0.071 0.090 0.353 0.152 0.070 0.645 0.396 0.170
Instructor-base w/o Inst. 0.246 0.263 0.282 0.024 0.036 0.102 0.301 0.059 0.036 0.584 0.415 0.134
w/ Inst. 0.201 0.239 0.254 0.023 0.038 0.096 0.164 0.052 0.069 0.582 0.401 0.127
Instructor-large w/o Inst. 0.248 0.295 0.332 0.037 0.051 0.125 0.264 0.068 0.016 0.710 0.546 0.146
w/ Inst. 0.234 0.266 0.286 0.031 0.044 0.108 0.226 0.062 0.034 0.672 0.527 0.141
@@
Cohere-Embed-v3 w/o Inst. 0.151 0.263 0.285 0.043 0.041 0.099 0.580 0.151 0.038 0.723 0.572 0.166
w/ Inst. 0.187 0.290 0.279 0.050 0.078 0.101 0.650 0.171 0.033 0.721 0.566 0.186
OpenAI-ada-002 w/o Inst. 0.256 0.293 0.310 0.031 0.058 0.133 0.197 0.100 0.042 0.732 0.834 0.167
w/ Inst. 0.106 0.248 0.239 0.026 0.058 0.118 0.114 0.096 0.048 0.673 0.824 0.137
OpenAI-3-small w/o Inst. 0.306 0.309 0.337 0.030 0.061 0.146 0.315 0.125 0.066 0.711 0.720 0.190
w/ Inst. 0.212 0.272 0.296 0.030 0.066 0.138 0.255 0.129 0.036 0.643 0.721 0.165
OpenAI-3-large w/o Inst. 0.373 0.341 0.420 0.034 0.101 0.240 0.291 0.164 0.075 0.901 0.896 0.236
w/ Inst. 0.342 0.314 0.375 0.050 0.136 0.212 0.339 0.211 0.074 0.877 0.894 0.250
@@
Latest instruction-aware decoder representational models (Wang et al., 2023a; Muennighoff et al., 2024) seem to start bridging this behavioral gap, pointing the promising direction of scaling decoder models for next-level representational abilities. Cohere’s proprietary Embed-English-v3.0 presents a strong gain brought by instructions. Considering its relatively early release time (Nov 02, 2023) compared to latest models that yield better absolute performance (OpenAI-v3, E5-Mistral, GritLM) on RAR-b, this behavior is surprising, and might be relevant to the efforts put in relevance feedback from LLMs  [^3] https://txt.cohere.com/introducing-embed-v3/. However, no public information is available whether Cohere’s current embedding model is an encoder or a decoder, and thus its strong instruction-following abilities can not add to our main narrative of the potentail for decoder representation models.
@@ 3.3 Multiple-choice Retrieval Setting
A special case in the models is Instructor, which is known to have seen the formats of the first 6 tasks through finetuning on Super-NI (Wang et al., 2022b), enabling it to achieve a better performance on these tasks compared to traditional S-DR models. Yet, its performance on the two settings of CSTS that we construct is stuck at chance level. Without solid control experiments, we hypothesize that this pattern stems from the subpar generalizablity of encoder models. Although the pattern can currently be empirically supported by the better performance achieved by similarly instruction-tuned decoder models including E5-Mistral, GritLM and OpenAI models, the generalization gap between encoders and decoders has not yet been systematically investigated on information retrieval, and we position this as a valuable research topic.
@@ 4.1 Behavioral Analysis through the lens of Temporal Reasoning
Lastly, for the Context+Fact setting, we further concatenate the context (the background knowledge of the entity in the question that is going to be asked) after the facts. On top of the Fact setting, the Context+Fact setting further reflects a model’s robustness towards irrelevant information (since the answer can be extracted only from facts, but not from contexts), and encoding capabilities for texts of varied length ranges (Xiao et al., 2023a). However, there potentially is unfair comparison in here because of the max sequence length each model is capable of encoding. For instance, OpenAI embedidng models are able to encode sequence up to 8192 tokens and are faced with more noise to be robust against. On the other hand, models with max sequence length of 512 need to deal with less noise. Therefore, the Context-Fact setting here is only for analysis purposes, and we do not average the scores from this setting into the average TR scores in the main table. For E5-Mistral and GritLM, we set its max sequence length to 8192 here to match with OpenAI models, for attaining a comparable understanding on the Context+Fact setting.
@@
Cohere-Embed-v3 w/o Inst. 0.015 0.019 0.359 0.197 0.085 0.275 0.162 0.159 0.151
@@ Appendix A Dataset Format
Dataset Instruction Query Doc
$\alpha$NLI Given the following start and end of a story, retrieve a possible reason that leads to the end. Start: Ron started his new job as a landscaper today. End: Ron is immediately fired for insubordination. Ron ignores his bosses’s orders and called him an idiot.
HellaSwag Given the following unfinished context, retrieve the most plausible ending to finish it. A man is sitting on a roof. He starts pulling up roofing on a roof.
PIQA Given the following goal, retrieve a possible solution. How do I ready a guinea pig cage for it’s new occupants? Provide the guinea pig with a cage full of a few inches of bedding made of ripped paper strips, you will also need to supply it with a water bottle and a food dish.
SocialiQA Given the following context and question, retrieve the correct answer. Context: Tracy didn’t go home that evening and resisted Riley’s attacks. Question: What does Tracy need to do before this? Find somewhere to go
@@
ARC-Challenge Retrieve the answer to the question. An astronomer observes that a planet rotates faster after a meteorite impact. Which is the most likely effect of this increase in rotation?” Planetary days will become shorter.
Winogrande Given the following sentence, retrieve an appropriate answer to fill in the missing underscored part. Sentence: Sarah was a much better surgeon than Maria so _ always got the easier cases.. Maria
TempReason-L1 Given the following question about time, retrieve the correct answer. What is the time 6 year and 4 month after Nov, 1185 Mar, 1192
TempReason-L2 pure Given the following question, retrieve the correct answer. Which employer did Jaroslav Pelikan work for in Jan, 1948? Valparaiso University
@@
TempReason-L3 pure Given the following question, retrieve the correct answer. Which employer did Jaroslav Pelikan work for before Concordia Seminary? Valparaiso University
@@
SpartQA Given the following spatial reasoning question, retrieve the right answer. We have three blocks. Lets call them A, B and C. Block B is below A. Block A is below C. Block A contains a medium yellow square. Block B has two medium blue squares. Medium blue square number one is touching the bottom edge of this block. Medium blue square number two is below a medium yellow square. Medium blue square number one is below the square which is below the medium yellow square. It is below the medium yellow square. Block C contains one medium black square. What is below the black thing? a medium yellow square that is in block A or a medium yellow square that is in block B? both of them
Math-pooled Retrieve the answer for the following math problem. {Omitted for brevity} {Omitted for brevity}
HumanEvalPack-MBPP Retrieve the answer for the following coding problem. Finish the following code based on the docstring: {Omitted for brevity} {Omitted for brevity}
CSTS-easy Retrieve an aspect and a sentence which are similar to the following sentence. 3 people wearing festive costumes and feathers dancing in a parade . In terms of ”The number of persons” retrieve a sentence similar to the following. Three uniquely dressed women dancing in a parade .
CSTS-hard Retrieve a condition under which the following two sentences are similar. Sentence1: 3 people wearing festive costumes and feathers dancing in a parade .; Sentence2: Three uniquely dressed women dancing in a parade . The number of persons.
@@ Appendix D Full setting Recall@10
Cohere-Embed-v3 w/o Inst. 0.223 0.400 0.423 0.073 0.082 0.180 0.886 0.275 0.080 0.790 0.739 0.266